Hi, I'm trying to find the answer for: "46 less than three times a number is equal to the number. What is the number?" I believe it is 23, or what is the number?
Answer by peekaboo0501(13)


46-3x = x
46= 4x
11.5 = x

46-3(11.5)=11.5
46-34.5=11.5
11.5=11.5